# Building a Macedonian Recipe Dataset: Collection, Parsing, and Comparative Analysis

### v1

@@ IV-B Ingredient Pair and Triplet Frequency
Ingredient Pair Count Ingredient Pair %
@@
Ingredient Triplet Count Ingredient Triplet %
@@
Recipe1M+ displays more balanced diversity in its top combinations. While baking staples remain prominent—butter & salt (9.2%), salt & sugar (8.8%), eggs & salt (7.7%)—the dataset lacks the extreme concentration around a single ingredient observed in the Macedonian collection. The appearance of olive oil & salt (5.3%) signals substantial Mediterranean and savory cooking content. The presence of baking powder in multiple top triplets reflects modern convenience baking, contrasting with the simpler flour-egg-sugar combinations typical of traditional Macedonian recipes.
A notable difference in co-occurrence density emerges from the raw counts. The Macedonian top pair appears in 17.4% of recipes, while Recipe1M+’s top pair appears in less than 1% of its collection. This disparity reflects not only dataset size differences but also greater recipe diversity in the international corpus. Recipe1M+ encompasses a broader range of culinary traditions, from Asian stir-fries to European sauces to American desserts, resulting in more fragmented co-occurrence patterns. The Macedonian dataset, by contrast, shows tighter clustering around culturally specific ingredient combinations.
@@ V-D Technical Parsing Challenges
Compound ingredients: Multi-word ingredient names (маслиново масло за прелив—"olive oil for dressing") sometimes confuse ingredient identity with preparation context, artificially fragmenting the ingredient space.

### v2

@@ IV-B Ingredient Pair and Triplet Frequency
Ingredient Pair % Ingredient Pair %
@@
Ingredient Triplet % Ingredient Triplet %
@@
Recipe1M+ displays relatively more balanced diversity compared to the Macedonian collection’s extreme concentration. While baking staples remain prominent, butter & salt (9.2%), salt & sugar (8.8%), eggs & salt (7.7%), the dataset lacks the extreme concentration around a single ingredient observed in the Macedonian collection. The appearance of olive oil & salt (5.3%) signals substantial Mediterranean and savory cooking content. The presence of baking powder in multiple top triplets reflects modern convenience baking, contrasting with the simpler flour-egg-sugar combinations typical of traditional Macedonian recipes.
A notable difference in co-occurrence density emerges from the raw counts. The Macedonian top pair appears in 17.4% of recipes, while Recipe1M+’s top pair appears 9.2% of its collection. This disparity reflects both dataset size differences and greater recipe diversity in the international corpus. Recipe1M+ encompasses a broader range of culinary traditions, from Asian stir-fries to European sauces to American desserts, resulting in more fragmented co-occurrence patterns. The Macedonian dataset, by contrast, shows tighter clustering around culturally specific ingredient combinations.
@@ V-D Technical Parsing Challenges
Compound ingredients: Multi-word ingredient names (маслиново масло за прелив (olive oil for dressing)) sometimes confuse ingredient identity with preparation context, artificially fragmenting the ingredient space.
@@ VI Conclusion
VII Code Availability
The code for dataset generation, parsing, and analysis is available at https://gitlab.finki.ukim.mk/ds4food/MKRecipes.